# Understanding Representation Dynamics of Diffusion Models via Low-Dimensional Modeling

### v2

@@ Appendix D Experimental Details
Each model is trained for 500 epochs using the Adam optimizer [Kingma, 2015] with a fixed learning rate of $1\times 10^{-4}$. The sliced Wasserstein distance is computed according to the implementation described in [Doan et al., 2024].
Experimental details for Table 1 and Table 2
For EDM, we use the official pre-trained checkpoints on ImageNet $64\times 64$ from [Karras et al., 2022], and for DiT, we use the released DiT-XL/2 model pre-trained on ImageNet $256\times 256$ from [Peebles and Xie, 2023]. As a baseline, we include the Hugging Face pre-trained MAE encoder (ViT-B/16) [He et al., 2022].
For diffusion models, features are extracted from the layer and timestep that achieve the highest probing accuracy, following [Xiang et al., 2023]. After feature extraction, we adopt the probing protocol from [Chen et al., 2024b], passing the extracted features through a probe consisting of a BatchNorm1d layer followed by a linear layer. To ensure fair comparisons, all input images are cropped or resized to $224\times 224$, matching the resolution used for MAE training.
For ensembling, we extract features from two additional timesteps on either side of the optimal timestep. Independent probes are trained on these timesteps, yielding five probes in total. At test time, we apply a soft-voting ensemble by averaging the output logits from all five probes for the final prediction. Specifically, let $\bm{W}_{t}\in\mathbb{R}^{K\times d}$ be the linear classifier trained on features from timestep $t$, and let $\bm{h}_{t}\in\mathbb{R}^{d}$ denote the feature representation of a sample at timestep $t$. Considering neighboring timesteps $t-2,t-1,t+1$, and $t+2$, our ensemble prediction is computed as: ^$\bm{\hat{}}{y}=\operatorname*{\text{arg~{}max}}\left(\frac{1}{5}\sum_{t=t-2}^{t+2}\bm{W}_{t}\bm{h}_{t}\right).$
We evaluate each method under varying levels of label noise, ranging from $0\%$ to $80\%$, by randomly mislabeling the specified percentage of training labels before applying linear probing. Performance is assessed on both the pre-training dataset and downstream transfer learning tasks. For pre-training evaluation, we use the images and classes from MiniImageNet [Vinyals et al., 2016] to reduce computational cost. For transfer learning, we evaluate on CIFAR100 [Krizhevsky et al., 2009], DTD [Cimpoi et al., 2014], and Flowers102 [Nilsback and Zisserman, 2008].
@@ Appendix E Proofs
We first provide some auxiliary results for proving Theorem 1.
@@ E.1 Groud truth posterior mean and optimal DAE
Suppose the data $\bm{x}_{0}$ is drawn from a noisy MoLRG data distribution with $K$-class and noise level $\delta$ introduced in 1. Then the optimal $\{\bm{U}\}$ minimizing the loss (2) is the ground truth basis defined in (3), and the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)$ admits the analytical form:
@@
By [Vincent, 2011] we see that the ground-truth score/posterior estimator calculated from the pdf is the global minimizer of the denoising score matching loss in (2), so here we first calculate the ground-truth score of the noisy MoLRG distribution and the corresponding posterior. We follow the same proof steps as in [Wang et al., 2024] Lemma $1$ with a change of variable. Let $\bm{c}_{k}=\begin{bmatrix}\bm{a}_{k}\\ \bm{e}_{k}\end{bmatrix}$ and $\widehat{\bm{U}}_{k}=\begin{bmatrix}\bm{U}_{k}^{\star}&\delta\widetilde{\bm{U}}_{k}^{\star}\end{bmatrix}$ where $d$ and $D$ denote the dimensions of signal space and noise space as in definition, we first compute the conditional pdf
@@
E.2 Proof of Theorem 1
We first state the formal version of Theorem 1.
To simplify the calculation of $\mathrm{SNR}$ as introduced in Section 3.3 on feature representations, which involves the expectation over the softmax term $w^{\star}_{k}$, we approximate $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ as follows:
@@ E.2 Proof of Theorem 1
In other words, we use $\hat{w}_{k}^{\star}$ in (10) to approximate $w^{\star}_{k}(\bm{x}_{t},t)$ in 1 by taking expectation inside the softmax with respect to $\bm{x}_{t}$. This allows us to treat $\hat{w}_{k}^{\star}$ as a constant when calculating $\mathrm{SNR}$, making the analysis more tractable while maintaining $\operatorname{\mathbb{E}}[\|\bm{U}_{l}^{\star}\hat{\bm{h}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)\|^{2}]\approx\operatorname{\mathbb{E}}[\|\bm{U}_{l}^{\star}\hat{\bm{h}}_{\textit{approx}}^{\star}(\bm{x}_{0},t)\|^{2}]$ for all $l\in[K]$. We verify the tightness of this approximation at Appendix C (Figure 14). With this approximation, we state the theorem as follows:
@@
Let data $\bm{x}_{0}$ be any arbitrary data point drawn from the MoLRG distribution defined in Assumption 1 and let $k$ denote the true class $\bm{x}_{0}$ belongs to. Then $\mathrm{SNR}$ introduced in Section 3.3 depends on the noise level $\sigma_{t}$ in the following form:
@@
Following the definition of $\mathrm{SNR}$ as defined in Section 3.3, Lemma 2 and the fact that $k\sim\text{Mult}(K,\pi_{k})$ with $\pi_{1}=\dots=\pi_{K}=1/K$, we can write
@@
That is, we consider a simplified form of the expected posterior mean from 1, where the expectation is taken inside the softmax argument (i.e., over $g_{k}^{\star}(\bm{x}_{t},t)$) to obtain tractable approximate weights $\hat{w}_{k}^{\star}$.
@@
$\bm{e}_{i}[a:b]\;\;\;\;\;\;\text{Slices of vector}\;\bm{e}_{i}\;\text{from}\;a\text{th}\;\text{entry to}\;b\text{th}\;\text{entry.}$.
@@
This, together with Lemma 1, yield (15).

### v3

@@ Appendix D Experimental Details
Each model is trained for 500 epochs using the Adam optimizer [kingma2014adam] with a fixed learning rate of $1\times 10^{-4}$. The sliced Wasserstein distance is computed according to the implementation described in [doan2024assessing].
Experimental details for Table˜1 and Table˜2
For EDM, we use the official pre-trained checkpoints on ImageNet $64\times 64$ from [karras2022elucidating], and for DiT, we use the released DiT-XL/2 model pre-trained on ImageNet $256\times 256$ from [peebles2023scalable]. As a baseline, we include the Hugging Face pre-trained MAE encoder (ViT-B/16) [he2022masked].
For diffusion models, features are extracted from the layer and timestep that achieve the highest probing accuracy, following [xiang2023denoising]. After feature extraction, we adopt the probing protocol from [chen2024deconstructing], passing the extracted features through a probe consisting of a BatchNorm1d layer followed by a linear layer. To ensure fair comparisons, all input images are cropped or resized to $224\times 224$, matching the resolution used for MAE training.
For ensembling, we extract features from two additional timesteps on either side of the optimal timestep. Independent probes are trained on these timesteps, yielding five probes in total. At test time, we apply a soft-voting ensemble by averaging the output logits from all five probes for the final prediction. Specifically, let $\bm{W}_{t}\in\mathbb{R}^{K\times d}$ be the linear classifier trained on features from timestep $t$, and let $\bm{h}_{t}\in\mathbb{R}^{d}$ denote the feature representation of a sample at timestep $t$. Considering neighboring timesteps t2,t1,t+1t-2,t-1,t+1, and $t+2$, our ensemble prediction is computed as: $\bm{\hat{}}{y}=\operatorname*{\text{arg~max}}\left(\frac{1}{5}\sum_{t=t-2}^{t+2}\bm{W}_{t}\bm{h}_{t}\right).$
We evaluate each method under varying levels of label noise, ranging from $0\%$ to $80\%$, by randomly mislabeling the specified percentage of training labels before applying linear probing. Performance is assessed on both the pre-training dataset and downstream transfer learning tasks. For pre-training evaluation, we use the images and classes from MiniImageNet [vinyals2016matching] to reduce computational cost. For transfer learning, we evaluate on CIFAR100 [krizhevsky2009learning], DTD [cimpoi14describing], and Flowers102 [nilsback2008automated].
@@ Appendix E Proofs
We first provide some auxiliary results for proving Theorem˜1.
@@ E.1 Groud truth posterior mean and optimal DAE
Suppose the data $\bm{x}_{0}$ is drawn from a noisy MoLRG data distribution with $K$-class and noise level $\delta$ introduced in ˜1. Then the optimal $\{\bm{U}\}$ minimizing the loss (2) is the ground truth basis defined in (3), and the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)$ admits the analytical form:
@@
By [vincent2011connection] we see that the ground-truth score/posterior estimator calculated from the pdf is the global minimizer of the denoising score matching loss in (2), so here we first calculate the ground-truth score of the noisy MoLRG distribution and the corresponding posterior. We follow the same proof steps as in [wang2024diffusion] Lemma $1$ with a change of variable. Let $\bm{c}_{k}=\begin{bmatrix}\bm{a}_{k}\\ \bm{e}_{k}\end{bmatrix}$ and $\widehat{\bm{U}}_{k}=\begin{bmatrix}\bm{U}_{k}^{\star}&\delta\widetilde{\bm{U}}_{k}^{\star}\end{bmatrix}$ where $d$ and $D$ denote the dimensions of signal space and noise space as in definition, we first compute the conditional pdf
@@
E.2 Proof of Theorem˜1
We first state the formal version of Theorem˜1.
To simplify the calculation of $\mathrm{SNR}$ as introduced in Section˜3.3 on feature representations, which involves the expectation over the softmax term $w^{\star}_{k}$, we approximate $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ as follows:
@@ E.2 Proof of Theorem˜1
In other words, we use $\hat{w}_{k}^{\star}$ in (10) to approximate $w^{\star}_{k}(\bm{x}_{t},t)$ in ˜1 by taking expectation inside the softmax with respect to $\bm{x}_{t}$. This allows us to treat $\hat{w}_{k}^{\star}$ as a constant when calculating $\mathrm{SNR}$, making the analysis more tractable while maintaining $\operatorname{\mathbb{E}}[\|\bm{U}_{l}^{\star}\hat{\bm{h}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)\|^{2}]\approx\operatorname{\mathbb{E}}[\|\bm{U}_{l}^{\star}\hat{\bm{h}}_{\textit{approx}}^{\star}(\bm{x}_{0},t)\|^{2}]$ for all $l\in[K]$. We verify the tightness of this approximation at Appendix˜C (Figure˜14). With this approximation, we state the theorem as follows:
@@
Let data $\bm{x}_{0}$ be any arbitrary data point drawn from the MoLRG distribution defined in Assumption 1 and let $k$ denote the true class $\bm{x}_{0}$ belongs to. Then $\mathrm{SNR}$ introduced in Section˜3.3 depends on the noise level $\sigma_{t}$ in the following form:
@@
Following the definition of $\mathrm{SNR}$ as defined in Section˜3.3, Lemma˜2 and the fact that $k\sim\text{Mult}(K,\pi_{k})$ with $\pi_{1}=\dots=\pi_{K}=1/K$, we can write
@@
That is, we consider a simplified form of the expected posterior mean from ˜1, where the expectation is taken inside the softmax argument (i.e., over $g_{k}^{\star}(\bm{x}_{t},t)$) to obtain tractable approximate weights $\hat{w}_{k}^{\star}$.
@@
$\bm{e}_{i}[a:b]\;\;\;\;\;\;\text{Slices of vector}\;\bm{e}_{i}\;\text{from}\;a\text{th}\;\text{entry to}\;b\text{th}\;\text{entry.}$
@@
This, together with Lemma˜1, yield (15).